OpenStudy (anonymous):

Help ME PLEASE!!!!!! Josiah is taking his friends to the movies. Each ticket costs $10, and popcorn is $5 a bag. There is a $3 service fee for the entire purchase. He has $75. If he buys 4 tickets, how many bags of popcorn can he buy? If x = number of tickets and y = number of bags of popcorn, which of the following inequalities could you use to answer the problem?

OpenStudy (aum):

Each ticket costs $10. He bought "x" tickets. Ticket cost = 10*x Each bag of popcorn costs $5. He bought "y" bags of popcorn. Popcorn cost = 5*y Service fee: $3 Total cost = 10x + 5y + 3 <= $75 is the inequality. 10x + 5y <= 72 If x = 4, then 40 + 5y <= 72 5y <= 32 y <= 6.4 So y = 6 bags of popcorn.

OpenStudy (aum):

He can buy a maximum of 6 bags of popcorn.

OpenStudy (anonymous):

thank you